Currently, the spot exchange rate is $0.85/A$, and the one-year forward exchange rate is $0.81/A$. One-year interest is 3.5% in the United States and 4.2% in Australia. You may borrow up to $1,000,000 or A$1,176,471, which is equivalent to $1,000,000 at the current spot rate.

  1. Determine if Interest Rate Parity (IRP) is holding between Australia and the United States.
  2. If IRP is not holding, explain in detail how you would realize certain profit in U.S. dollar terms.
  3. Explain how IRP will be restored as a result of arbitrage transactions you carry out above.
Answer #1

Ans Current spot Exchange Rate = 0.85 /A$ One year four ward Exchange Rate = $0.81/A Interest lade in US = 3.5% Interest RateBarrow 1176471 in Australian market @of 4.2% Convort this amount into USD @ 0.85/AS = 1176471 x 0.85 O = $ 1000000 Invest thi